OpenStudy (elise_a18):

What is 3 to the power of 2 over 3 equal to? cube root of 9 square root of 9 cube root of 27 square root of 27

OpenStudy (elise_a18):

@confluxepic @Coolguy25 @jim_thompson5910 @sleepyjess @aaronq @SamsungFanBoy @linn99123 @Omgthatgirl @pooja195

OpenStudy (anonymous):

27 if u meant 9^(3/2) when we do fractional indices we root the base by the denominator then exponent it to the numerator so for 9^(3/2) we do square root 9 =3 then we cube it 3^3=27 or if u meant (9^3)/2 the answer is 729/2=364.5

OpenStudy (elise_a18):

what is \[3\frac{ 2 }{ 3 }\] equal to? \[\sqrt[3]{9}\] \[\sqrt[2]{9}\] \[\sqrt[3]{27}\] or\[\sqrt[2]{27}\]

OpenStudy (bibby):

fractional exponents are roots. \(\huge a^{\frac{1}{x}}=\sqrt[x]a\\(3)^\dfrac{2}{3}=\sqrt[3]{3^2}\)

OpenStudy (anonymous):

lol it was a joke

OpenStudy (anonymous):

got a lot of people needing my help sorry

OpenStudy (elise_a18):

so what would it be bibby? i dont know how to find the answer with that info

OpenStudy (bibby):

the answer is \(\huge \sqrt[3]{3^2}\). simplify it yourself.

OpenStudy (elise_a18):

that's the problem. i dont know how :c

OpenStudy (bibby):

what is 3^2?

OpenStudy (elise_a18):

9

OpenStudy (elise_a18):

oh wow

OpenStudy (bibby):

I know right

OpenStudy (bibby):

:p

OpenStudy (elise_a18):

Sorry. It's 2am where I am XP

OpenStudy (elise_a18):

XI

OpenStudy (bibby):

UK?

OpenStudy (elise_a18):

I'm not telling you where I live!

OpenStudy (elise_a18):

O-O